# Learning with pyCub: A Simulation and Exercise Framework for Humanoid Robotics

### v2

@@ 1 Introduction
There are a number of good robotics textbooks, such as [25] or [13] with a great video supplement, but few of them offer hands-on exercises, toolboxes, or robot simulators. The book of Corell et al. [3] is accompanied by a GitHub repository with Matlab and Mathematica code. Peter Corke developed a Python Robotics Toolbox [2]. This toolbox provides tools for both dynamics and kinematics of robots and is a successor to the Robotic Toolbox written in MATLAB [1]. Tutorials on robot differential kinematics using the Python Robotics Toolbox were presented by Haviland and Corke [10, 9]. et al.Sakai [24] proposed an educational textbook with Python examples to explain several robotic algorithms for autonomous navigation. One part of our work is similar in the sense that we provide exercises that should teach core robotics.
@@
Our presented framework was created as a tool for our university course. Our course is on humanoid robots, but due to the universal nature of such a platform, it can be used to teach several robotic topics, from kinematics and dynamics, to walking, gaze control, grasping, and human-robot interaction. Our objective thus was to create a robot simulator that is easily accessible to students and can be used for hands-on exercises of the course. We selected iCub [16] as it is a full humanoid and an open-source research platform. The dimensions of the robot resemble a four-year-old child, and the geometry and kinematics are designed to mimic the human body as much as possible (at the joint level, not at the muscle level). There are 53 active Degree of Freedoms. A unique feature of iCub is the whole-body artificial skin with more than 4000 individual tactile sensors. The robot has a legacy simulator “iCub SIM” [26], a newer simulator in Gazebo [17] and offers its own training course [20]. However, getting the environment running is an intricate process, and the default control is in C++ using a Yet Another Robot Platform (YARP[15] middleware, which is not suitable for beginners—there are optional Python bindings available, but as these are created from the C++ code, the syntax can be hard to grasp for less experienced users. After running two editions of the course using the iCub Gazebo, even with Docker containers and other infrastructure prepared, we have evaluated the feedback forms and realized that the students complained about the complexity of the whole system. Thus, we decided to simplify the whole pipeline and create a simulator using Python, which is a more user-friendly language. The comparison of existing iCub simulators is in Table 1—the original simulators and a neuromorphic (and event-based) simulator [6]. We are not aiming to replace the existing simulators, but to provide a more accessible way to iCub and humanoid robotics in general.
Table 1: Comparison of pyCub with existing iCub simulators (iCub SIM [26], iCub Gazebo [17], and neuromorphic iCub [6]). *Skin to Gazebo was added later by the community. **The planners are coming from the connection with YARP. ***By default, can be changed.****Event-driven skin.
@@
YARP Yes Yes* Yes** DART*** OGRE
Neuromorphic
iCub
Python None No Yes**** No MuJoCo OpenGL
@@
We present pyCub, a Python physics-based simulator of the humanoid robot iCub. The simulator contains all limbs of the robot, together with the possibility to visualize view from the robot’s cameras, and with simulated artificial whole-body skin consisting of more than 4000 individual tactile sensors. The framework provides easy-to-use functions to control the robot in joint (position and velocity) and Cartesian space, and effortless access to proprioceptive, visual, and tactile sensors of the robot. Everything is written and accessible in Python, allowing even less advanced users to operate. The environment is fully customizable and allows to insert other objects—see Fig. 1 for an example. In addition to the simulator, we provide exercises of different difficulty levels ranging from basic control and movements to tasks typical for humanoid robots such as gazing, grasping, or whole-body skin processing and obstacle avoidance. The exercises can be further customized and allow difficulty adjustments and thus are suitable for learning (humanoid) robotics on different difficulty levels. The framework was deployed in two runs of a course on humanoid robotics.
@@ 2 Simulation Environment
2.1 Technical Details
The physics-backbone of the simulator is realized with the PyBullet engine [4]. This module was selected because it is open-source, multi-platform and does not require excessive computational resources—mainly compared to recent physical simulators such as Orbit (NVIDIA Issac Sim) [18]. The schematic of the simulator package is shown in Fig. 2.
@@ 2.1.1 Time Management
2.1.2 GUI
The visualization part is achieved with the Open3D library [29]. PyBullet has its own built-in GUI, but it is not customizable, slow, and does not support custom simulation stepping. Therefore, we developed a new GUI that removes all the drawbacks. In addition, it allows one to easily visualize things that are not part of the simulation, e.g., visualization of gaze vectors—see Fig. 3. In native mode, GUI runs as a standalone window. However, for compatibility reasons, a web-based version can be turned on. Both versions have the possibility to view output from the eyes and allow to save RGB and depth images to file.
@@ 2.1.2 GUI
2.1.3 Customizations
The environment is fully customizable through .yaml configuration files. The files are loaded when initializing an instance of the simulator, and thus users can create different configurations for different exercises/tasks/demos. The basic options allow to enable/disable GUI, artificial skin, eye-cameras output, logging, or self collisions. Further, the configurations allow us to set an initial position for every joint of the robot, the end-effector used for Cartesian control, or simulation step size. Finally, users can insert additional objects into the simulation. The configurations support Unified Robot Description Format (URDF) files, or directly .obj triangular-mesh files. One can further specify the initial position, color, and whether the object is subject to gravity and can be moved (e.g, ball) or not (e.g., table).
2.1.4 Installation
To provide easy installation, the framework can be installed from PyPI. Alternatively, we provide several versions of prebuilt Docker images or a version runnable entirely online using the GitPod [^1] https://www.gitpod.io/ platform. Everything is available at https://rustlluk.github.io/pyCub.
@@ 2.2 Control Modes
Cartesian Position Control The Cartesian mode allows to move the selected end-effector to a given 6D pose (3D position and 3D orientation). Internally, the function computes Inverse Kinematics (IK) for the given pose and calls position control with joint positions computed using IK—the IK in PyBullet uses Damped Least Squares method with null-space support. There is no underlying planner, i.e., there are no checks whether the pose is reachable or whether the path is collision free. However, it creates a great platform for learning how movements without planner or path smoothing look.
@@ 2.3 Artificial Skin
The artificial skin is an integral component of the robot. However, the real robot contains more than 4000 individual sensors, making the simulation challenging. The sensors are also not binary, i.e., they can return value with 8-bit resolution. We decided to simulate the skin using raycasting. From every skin sensor (taxel) we cast a short ray with a given maximum length. If the ray hits something (other part of the robot or object in the environment), the output of the sensor is computed proportionally to the length of the ray. Using this method, we are able to simulate the intensity of contact and also the fact that for the real skin, even taxels around the touch point are lightly activated because of the deformation of the skin. The raycasting needs to be done in every step of the simulation, and, even though the raycasting is optimized and done in batch, it is still computationally demanding. Thus, we decided to first check (for every skin part) whether there is a close object (using overlaps of bounding boxes of objects) and perform the raycasting only if necessary. The values for each activated taxels are saved in a dictionary inside the simulator and binary activations are shown as green points in the visualization—see Fig. 4.
@@ 3 Exercises
This Section introduces and describes the exercises available in the framework. We prepared several exercises that, in our opinion, span important tasks in robotics. We will describe those in recommended order based on their difficulty. However, the exercises can be further adjusted to change the difficulty level. The exercises always contain a configuration file, an assignment script, and a testing script. Videos of each exercise are available at https://rustlluk.github.io/pyCub and in the accompanying video.
@@ 3.1 Push the Ball!
The first exercise is purely for the purpose of learning how to control the simulator. The goal is to get a ball as far from a table as possible—see Fig. 1 for visualization. Basically, every imaginable approach is allowed. Whether hitting, pushing, or grasping and throwing the ball. Upon completion, users should be able to control the robot in some of the modes and understand basic functionalities. During this exercise, the self-collisions are turned off—it is visible in some examples in the accompanying video.
The testing script for this exercise computes the distance of a ball and a table after two seconds after the last movement of the robot and provides the distance in meters.
@@ 3.2 Smooth Movements
As described in Section 2.2, the Cartesian interface of the simulator does not include any underlying planner. This exercise aims to provide insight on how to move in a straight line or how to “draw” a shape, e.g., a circle. In addition, the movements should be “smooth”, i.e., a continuous movement without any jerks.
@@ 3.3 Gaze
Gaze control is a key ability so that the robot can bring some objects into the field of view. The goal of gaze can be described as focusing on the object of interest using, mainly, joints in the head and eyes. There is a gaze control implemented in the real iCub robot [22] that is capable of fixing to a given 3D using all joints in the neck and eyes. In this exercise, the problem is simplified to fixating the view of the robot using neck joints to a ball moving in XY-plane.
In the default version of the exercise, the users are given two vectors: (i) vector where the robot is looking right now; (ii) vector from the robot to a ball, i.e., where the robot should be looking. See Fig. 3 for visualization. The task is then to use these two vectors to fixate to the ball.
The two vectors should lead users to use a fundamental algebra to find an angle $\theta$ between two vectors, for example, as:
@@
3.4 Reactive Control and Skin Processing
@@
The exercise has different levels of difficulty depending on the limb in contact and the number of simultaneous contacts. The testing script here runs the different levels of difficulty. At the moment, there is no automatic evaluation and performance must be checked manually.
@@ 3.5 Grasping
Another important skill for humanoids is object handover between the human and the robot. The robot must be reactive (the previous exercise) and must be able to grasp objects. Robotic grasping has changed a lot during its development. From grasping based on friction cone and wrenches, through physics-based grasping to Machine Learning (ML) and Deep Learning (DL)-based approaches [5].
@@ 4 Conclusion, Discussion, and Future Work
We presented pyCub, an unified educational platform consisting of a physics-based simulation of the humanoid robot iCub and exercises that aim to teach core robotic tasks and skills. The whole framework is open-source and available at https://rustlluk.github.io/pyCub. The exercises cover basic robot control in joint and Cartesian space, more complicated Cartesian movements, gazing at a moving target, reactive movements in velocity space to avoid obstacles detected using signal processing from whole-body artificial skin, and grasping based on computer vision techniques.
The simulation and exercises were deployed in two runs of a course on humanoid robotics. Previous runs were taught using iCub’s original Gazebo simulator controlled through a middleware YARP [15] using C++. Switching to pyCub, a Python-based simulator changed the dynamics of the class. Although the students were engineering students with experience with C++, the whole system was troublesome to learn and most of the complaints in the post-semester survey were about YARP and C++. Using pyCub allowed them to focus more on the robotic aspects of the exercises rather than on new middleware or compilation issues. For example, although the first exercise is mainly intended to get used to the simulator, the students showed that it is also an engaging exercise for them. The goal is to get the ball as far as possible from a table, and the expected solution is to simply hit the ball with an arm. However, there were many solutions such as grasping and throwing or even kicking the ball—see the accompanying video or the website of the project at https://rustlluk.github.io/pyCub.

### v3

@@ 1 Introduction
There are several good robotics textbooks, such as [26] or [14] with a great video supplement, but few of them offer hands-on exercises, toolboxes, or robot simulators. The book of Corell et al. [4] is accompanied by a GitHub repository with Matlab and Mathematica code. Peter Corke developed a Python Robotics Toolbox [3]. This toolbox provides tools for both dynamics and kinematics of robots and is a successor to the Robotic Toolbox written in MATLAB [2]. Tutorials on robot differential kinematics using the Python Robotics Toolbox were presented by Haviland and Corke [11, 10]. et al.Sakai [25] proposed an educational textbook with Python examples to explain several robotic algorithms for autonomous navigation. One part of our work is similar in the sense that we provide exercises that should teach core robotics.
@@
Our presented framework was created as a tool for our university course. Our course is on humanoid robots, but due to the universal nature of such a platform, it can be used to teach several robotic topics, from kinematics and dynamics, to walking, gaze control, grasping, and human-robot interaction. Our objective thus was to create a robot simulator that is easily accessible to students and can be used for hands-on exercises in the course. We selected iCub [17] as it is a full humanoid and an open-source research platform. The dimensions of the robot resemble a four-year-old child, and the geometry and kinematics are designed to mimic the human body as much as possible (at the joint level, not at the muscle level). There are 53 active Degrees of Freedom (DoF). A unique feature of iCub is the whole-body artificial skin with more than 4000 individual tactile sensors. The robot has a legacy simulator “iCub SIM” [27], a newer simulator in Gazebo [18] and offers its own training course [21]. However, getting the environment running is an intricate process, and the default control is in C++ using the Yet Another Robot Platform (YARP[16] middleware, which is not suitable for beginners—there are optional Python bindings available, but as these are created from the C++ code, the syntax can be hard to grasp for less experienced users—others tackled this by creating their own wrappers, e.g., [1]. After running two editions of the course using the iCub Gazebo, even with Docker containers and other infrastructure prepared, we have evaluated the feedback forms and realized that the students complained about the complexity of the whole system. Thus, we decided to simplify the whole pipeline and create a simulator using Python, which is a more user-friendly language. The comparison of existing iCub simulators is in Table 1—the original simulators and a neuromorphic (and event-based) simulator [7]. We are not aiming to replace the existing simulators, but to provide a more accessible way to iCub and humanoid robotics in general.
Table 1: Comparison of pyCub with existing iCub simulators (iCub SIM [27], iCub Gazebo [18], and neuromorphic iCub [7]).
@@
YARP Yes Yesa Yesb DARTc OGRE
Neuromorphic iCub Python None No Yes No MuJoCo OpenGL
@@
a Skin to Gazebo was added later by the community.
b The planners are coming from the connection with YARP.
c By default, can be changed.
We present pyCub, a Python physics-based simulator of the humanoid robot iCub. The simulator contains all limbs of the robot, together with the possibility to visualize the view from the robot’s cameras, and with simulated artificial whole-body skin consisting of more than 4000 individual tactile sensors. The framework provides easy-to-use functions to control the robot in joint (position and velocity) and Cartesian space, and effortless access to proprioceptive, visual, and tactile sensors of the robot. Everything is written and accessible in Python, allowing even less advanced users to operate. The environment is fully customizable and allows to insert other objects—see Fig. 1 for an example. In addition to the simulator, we provide exercises of different difficulty levels ranging from basic control and movements to tasks typical for humanoid robots such as gazing, grasping, or whole-body skin processing and obstacle avoidance. The exercises can be further customized and allow difficulty adjustments and thus are suitable for learning (humanoid) robotics on different difficulty levels. The framework was deployed in two runs of a course on humanoid robotics.
@@ 2 Simulation Environment
2.1 Technical Details.
The physics-backbone of the simulator is realized with the PyBullet engine [5]. This module was selected because it is open-source, multi-platform and does not require excessive computational resources—mainly compared to recent physical simulators such as Orbit (NVIDIA Issac Sim) [19]. The schematic of the simulator package is shown in Fig. 2. The full documentation (and API) can be found at https://rustlluk.github.io/pyCub/documentation/.
@@
2.1.2 GUI.
The visualization part is achieved with the Open3D library [30]. PyBullet has its own built-in GUI, but it is not customizable, slow, and does not support custom simulation stepping. Therefore, we developed a new GUI that removes all the drawbacks. In addition, it allows one to easily visualize things that are not part of the simulation, e.g., visualization of gaze vectors—see Fig. 3. In native mode, GUI runs as a standalone window. However, for compatibility reasons, a web-based version can be turned on. Both versions have the possibility to view output from the eyes and allow to save RGB and depth images to a file.
2.1.3 Customizations.
The environment is fully customizable through .yaml configuration files. The files are loaded when initializing an instance of the simulator, and thus users can create different configurations for different exercises/tasks/demos. The basic options allow to enable/disable GUI, artificial skin, eye-cameras output, logging, or self collisions. Further, the configurations allow us to set an initial position for every joint of the robot, the end-effector used for Cartesian control, or the simulation step size. Finally, users can insert additional objects into the simulation. The configurations support Unified Robot Description Format (URDF) files, or directly .obj triangular-mesh files. One can further specify the initial position, color, and whether the object is subject to gravity and can be moved (e.g, ball) or not (e.g., table).
2.1.4 Installation.
To provide easy installation, the framework can be installed from PyPI. Alternatively, we provide several versions of prebuilt Docker images or a version runnable entirely online using the GitPod [^1] https://www.gitpod.io/ platform. Everything is available at https://rustlluk.github.io/pyCub. The platform was tested on several Linux distributions, Windows 10 and 11, and several versions of macOS—the ease of installation differ between systems and given computers.
@@ 2.2 Control Modes
Cartesian Position Control. The Cartesian mode allows to move the selected end-effector to a given 6D pose (3D position and 3D orientation). Internally, the function computes Inverse Kinematics (IK) for the given pose and calls position control with joint positions computed using IK—the IK in PyBullet uses Damped Least Squares method with null-space support. There is no underlying planner, i.e., there are no checks whether the pose is reachable or whether the path is collision free. However, it creates a great platform for learning how movements without a planner or path smoothing look.
@@ 2.3 Artificial Skin
The artificial skin is an integral component of the robot. However, the real robot contains more than 4000 individual sensors, making the simulation challenging. The sensors are also not binary, i.e., they can return a value with 8-bit resolution. We decided to simulate the skin using raycasting. From every skin sensor (taxel) we cast a short ray with a given maximum length. If the ray hits something (other part of the robot or object in the environment), the output of the sensor is computed proportionally to the length of the ray. Using this method, we are able to simulate the intensity of contact and also the fact that, for the real skin, even taxels around the touch point are lightly activated because of the deformation of the skin. The raycasting needs to be done in every step of the simulation, and, even though the raycasting is optimized and done in batch, it is still computationally demanding. Thus, we decided to first check (for every skin part) whether there is a close object (using overlaps of bounding boxes of objects) and perform the raycasting only if necessary. The values for each activated taxels are saved in a dictionary inside the simulator and binary activations are shown as green points in the visualization—see Fig. 4.
@@ 3 Exercises
This Section introduces and describes the exercises available in the framework. We prepared several exercises that, in our opinion, span important tasks in robotics. We will describe those in recommended order based on their difficulty. However, the exercises can be further adjusted to change the difficulty level. The exercises always contain a configuration file, an assignment script, and a testing script. Videos of each exercise are available at https://rustlluk.github.io/pyCub.
@@ 3.1 Push the Ball!
The first exercise is purely for the purpose of learning how to control the simulator. The goal is to get a ball as far from a table as possible—see Fig. 1 for visualization. Basically, every imaginable approach is allowed. Whether hitting, pushing, or grasping and throwing the ball. Upon completion, users should be able to control the robot in some of the modes and understand basic functionalities.
The testing script for this exercise computes the distance between a ball and a table after two seconds after the last movement of the robot and provides the distance in meters.
@@ 3.2 Smooth Movements
As described in Section 2.2, the Cartesian interface of the simulator does not include any underlying planner. This exercise aims to provide insight into how to move in a straight line or how to “draw” a shape, e.g., a circle. In addition, the movements should be “smooth”, i.e., a continuous movement without any jerks.
@@ 3.3 Gaze
Gaze control is a key ability so that the robot can bring some objects into the field of view. The goal of gaze can be described as focusing on the object of interest using, mainly, joints in the head and eyes. There is a gaze control implemented in the real iCub robot [23] that is capable of fixing to a given 3D point using all joints in the neck and eyes. In this exercise, the problem is simplified to fixating the view of the robot using neck joints to a ball moving in XY-plane.
In the default version of the exercise, the users are given two vectors: (i) vector where the robot is looking right now; (ii) vector from the robot to a ball, i.e., where the robot should be looking. See Fig. 3 for visualization. The task is then to use these two vectors to fixate on the ball.
The two vectors should lead users to use fundamental algebra to find an angle $\theta$ between two vectors, for example, as:
@@
3.4 Reactive Control and Touch Processing
@@
The exercise has different levels of difficulty depending on the limb in contact and the number of simultaneous contacts. The testing script here runs the different levels of difficulty. At the moment, there is no automatic evaluation, and performance must be checked manually.
@@ 3.5 Grasping
Robotic grasping has changed a lot during its development. From grasping based on friction cone and wrenches, through physics-based grasping to Machine Learning (ML) and Deep Learning (DL)-based approaches [6].
@@ 4 Conclusion, Discussion, and Future Work
We presented pyCub, a unified educational platform consisting of a physics-based simulation of the humanoid robot iCub and exercises that aim to teach core robotic tasks and skills. The whole framework is open-source and is available at https://rustlluk.github.io/pyCub. The exercises cover basic robot control in joint and Cartesian space, more complicated Cartesian movements, gazing at a moving target, reactive movements in velocity space to avoid obstacles detected using signal processing from whole-body artificial skin, and grasping based on computer vision techniques.
The simulation and exercises were deployed in two runs of a course on humanoid robotics. Previous runs were taught using iCub’s original Gazebo simulator controlled through a middleware YARP [16] using C++. Switching to pyCub, a Python-based simulator changed the dynamics of the class. Although the students were engineering students with experience with C++, the whole system was troublesome to learn and most of the complaints in the post-semester survey were about YARP and C++. Using pyCub allowed them to focus more on the robotic aspects of the exercises rather than on new middleware or compilation issues. For example, although the first exercise is mainly intended to get used to the simulator, the students showed that it is also an engaging exercise for them. The goal is to get the ball as far as possible from a table, and the expected solution is to simply hit the ball with an arm. However, there were many solutions such as grasping and throwing or even kicking the ball—see the accompanying video at https://rustlluk.github.io/pyCub.